# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ 3. Preliminaries
An embedding is an injective homomorphism $h\colon\mathfrak{A}\rightarrow\mathfrak{B}$ that additionally satisfies the following condition: for every $k$-ary relation symbol $R\in\tau$ and $\bar{t}\in A^{k}$ we have $h(\bar{t})\in R^{\mathfrak{B}}$ only if $\bar{t}\in R^{\mathfrak{A}}.$ An isomorphism is a surjective embedding. Two structures $\mathfrak{A}$ and $\mathfrak{B}$ are isomorphic if there exists an isomorphism from $\mathfrak{A}$ to $\mathfrak{B}$. An automorphism is an isomorphism from $\mathfrak{A}$ to $\mathfrak{A}$; we denote by $\mathrm{Aut}(\mathfrak{A})$ the set (group) of all automorphisms of $\mathfrak{A}$. The orbit of a tuple $\bar{t}\in A^{k}$ in $\mathfrak{A}$ is the set $\{g(\bar{t})\mid g\in\mathrm{Aut}(\mathfrak{A})\}$. Any tuples belonging to the same orbit satisfy precisely the same first-order formulas over $\mathfrak{A}$. The orbit equivalence relation provides a natural way to factorize relational structures; in the present paper, we will only need the following specific case. Given a relational structure $\mathfrak{A}$ and a subgroup $G\subseteq\mathrm{Aut}(\mathfrak{A})$, we denote by $\mathfrak{A}_{\!{}_{\vphantom{/}\smash{/G}}}$ the structure $\mathfrak{A}/E$ for $E\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{(x,y)\in A^{2}\mid\exists g\in G\colon g(x)=y\}$. As mentioned in the introduction, a polymorphism of $\mathfrak{A}$ is a homomorphism from a finite power of $\mathfrak{A}$ into $\mathfrak{A}$.
Universal algebra.
@@
An (equational) condition is a set of identities, i.e. formal expressions of the form $s\approx t$, where $s$ and $t$ are terms over a common set of function symbols. An equational condition is height-1 if it contains neither nested terms nor terms consisting of a single variable. We say that $\mathrm{Pol}(\mathfrak{A})$ (or some set of operations on $A$) satisfies an equational condition $\Sigma$ if the function symbols can be interpreted as elements of $\mathrm{Pol}(\mathfrak{A})$ so that, for each identity $s\approx t$ in $\Sigma$, the equality $s=t$ holds for any evaluation of variables in $A$. An example is the cyclic identity of arity $n$ given by $f(x_{1},\dots,x_{n})\approx f(x_{2},\dots,x_{n},x_{1})$, for a symbol $f$ of arity $n$. We say that an operation is cyclic if it satisfies the cyclic identity.
@@ 4.1. Injectivity and Datalog reductions
In the context of Conjecture 1.4, the concepts of CSP-injectivity and having no algebraicity are closely related. On the one hand, if $\mathfrak{B}$ is a homogeneous structure with no algebraicity and its relations only contain tuples with pairwise distinct entries, then it is CSP-injective [16, Lemma 4.3.6]. On the other hand, we can prove the following.
@@
Pol-injectivity could be viewed as a universal-algebraic counterpart of CSP-injectivity, despite the fact that the two properties cannot hold simultaneously in $\omega$-categorical structures.
@@
The two notions are connected in the sense that from CSP-injective templates one can produce a Datalog-interreducible template by adding the relation $I_{4}$ from Example 2.3.
@@
The expansion of any CSP-injective structure $\mathfrak{A}$ by the relation $I_{4}$ is a Pol-injective structure $\mathfrak{A}_{I_{4}}$, and $\mathrm{CSP}(\mathfrak{A})$ and $\mathrm{CSP}(\mathfrak{A}_{I_{4}})$ are Datalog-interreducible.
Datalog reductions between CSPs are typically specified using factorization-free Datalog-interpretations with parameters [1, 39], which are a particular case of logical interpretations with parameters. In the present paper, we only need a very special case of Datalog-reductions, which can be easily explained on an informal level.
@@
Datalog is defined by adding formation rules to the existential positive fragment of first-order logic whose semantics is defined with inflationary fixed-points of arbitrary operators. Every existential positive first-order formula is a Datalog formula and, if $\phi(\bar{x})$ is a Datalog formula over some relational signature $\tau\cup\{R\}$ with $R\notin\tau$, then $[\mathsf{ifp}_{R}\,\phi](\bar{x})$ is a Datalog formula over the signature $\tau$ whose semantics is given as follows. For a $\tau$-structure $\mathfrak{A}$ and a tuple $\bar{a}$ over $A$ matching the arity of $\bar{x}$, say $k$, we have $\mathfrak{A}\models[\mathsf{ifp}_{R}\,\phi](\bar{a})$ if and only if $\bar{a}$ is contained in the inflationary fixed-point of the operator $F_{\phi,R}(X)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{\bar{a}\in A^{k}\mid\bar{a}\in X\text{ or }(\mathfrak{A};X)\models\phi(\bar{a})\}$, i.e., in the limit of the sequence $\emptyset$, $F_{\phi,R}(\emptyset)$, $F_{\phi,R}^{2}(\emptyset)$, $\dots$ For example, the Datalog formula $[\mathsf{ifp}_{T}\,(x<z)\vee(\exists y\ldotp T(x,y)\wedge T(y,z))](x,z)$ computes the transitive closure of $<$.
@@
For a relational $\tau$-structure $\mathfrak{A}$, we say that $\mathrm{CSP}(\mathfrak{A})$ is solvable in Datalog if there exists a Datalog sentence that defines the complement of $\mathrm{CSP}(\mathfrak{A})$, i.e. the class of all finite $\tau$-structures which do not homomorphically map to $\mathfrak{A}$. For example, $\mathrm{CSP}(\mathbb{Q};<)$ is solvable in Datalog using the sentence $\exists u[\mathsf{ifp}_{T}\,(x<z)\vee(\exists y\ldotp T(x,y)\wedge T(y,z))](u,u)$. It is not hard to see that every Datalog formula is specified by a finite set of rules of the form $R(\bar{x})\Leftarrow R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{m}(\bar{x}_{m})$ where $R$ is a fixed-point variable. In the case of the Datalog formula from above computing the transitive closure of $<$, for example, the rules are $T(x,z)\Leftarrow(x<z)$ and $T(x,z)\Leftarrow T(x,y)\wedge T(y,z)$.
@@
As mentioned above, when it comes to Datalog-reductions between CSPs, we will only need the following simplified version. For a $\tau$-structure $\mathfrak{A}$ and $\sigma$-structure $\mathfrak{A^{\prime}}$, we say that $\mathrm{CSP}(\mathfrak{A^{\prime}})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$ if there exists a mapping $\mathcal{I}$ from finite $\sigma$-structures to finite $\tau$-structures, where $\mathcal{I}(\mathfrak{J})$ can be defined from $\mathfrak{J}$ using Datalog formulas, so that $\mathfrak{J}\rightarrow\mathfrak{A^{\prime}}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow\mathfrak{A}$. It is not hard to see that Datalog reductions compose, i.e., if $\mathrm{CSP}(\mathfrak{A}^{\prime\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A}^{\prime})$ and $\mathrm{CSP}(\mathfrak{A}^{\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$, then $\mathrm{CSP}(\mathfrak{A}^{\prime\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$.
@@ 4.2. Wreath products
Let $\mathfrak{A}$ and $\mathfrak{B}$ be structures over disjoint relational signatures $\tau$ and $\sigma$. Moreover, let $E$ be a fresh binary symbol. The wreath product $\mathfrak{A}\wr\mathfrak{B}$ is the structure over the signature $\tau\cup\{E\}\cup\sigma$ with domain $A\times B$ whose relations are defined as follows. For every $R\in\tau$ of arity $k$:
$R^{\mathfrak{A}\wr\mathfrak{B}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=$$\{((x_{1},y_{1}),\dots,(x_{k},y_{k}))\in(A\times B)^{k}\mid(x_{1},\dots,x_{k})\in R^{\mathfrak{A}}\text{ and }y_{1}=\cdots=y_{n}\},$
$E^{\mathfrak{A}\wr\mathfrak{B}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=$$\{((x_{1},y_{1}),(x_{2},y_{2}))\in(A\times B)^{2}\mid y_{1}=y_{2}\},$
and for every $S\in\sigma$ of arity $m$:
$S^{\mathfrak{A}\wr\mathfrak{B}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{((x_{1},y_{1}),\dots,(x_{m},y_{m}))\in(A\times B)^{m}\mid(y_{1},\dots,y_{m})\in S^{\mathfrak{B}}\}.\hskip 71.13188pt$
@@
Given groups $G$ and $H$ acting on sets $A$ and $B$, respectively, their wreath product $G\wr H$ is given by $G^{B}\times H$ acting on $A\times B$ via $\left((g_{b})_{b\in B},h)\right)(a,b)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(g_{h(b)}(a),h(b)).$ The wreath product $G\wr H$ again forms a group, with the neutral element $((e_{G})_{b\in B},e_{H})$ and the group operation $((g_{b})_{b\in B},h))\cdot((g^{\prime}_{b})_{b\in B},h^{\prime}))\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=((g_{b}\cdot g^{\prime}_{h^{-1}(b)})_{b\in B},h\cdot h^{\prime}))$. Note that the permutation group on $A\times B$ induced by the component-wise action of the Cartesian product $G\times H$ is contained by that induced by the action of $G\wr H$ thereon: the permutations induced stemming from the former are represented in $G\wr H$ by all elements of the form $((g)_{b\in B},h)$.
@@
4.3. Putting everything together
@@
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(\mathbb{Q};<)\wr\mathfrak{B}$. By Example 1.3 and Example 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. We then define $(\mathbb{Q};<)^{\mathfrak{B}}$ as the expansion of $\mathfrak{B}^{\prime}$ by the binary inequality $\neq$. Note that $(\mathbb{Q};<)^{\mathfrak{B}}$ has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $(\mathbb{Q};<)^{\mathfrak{B}}_{I_{4}}$. We obtain the structure $\mathbb{Q}^{\mathfrak{A}}$ as a reduct of $(\mathbb{Q};<)^{\mathfrak{B}}$ by removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$; but note that $\mathbb{Q}^{\mathfrak{A}}$ can also be defined directly from $\mathfrak{A}$.
@@
We claim that $\mathbb{Q}^{\mathfrak{A}}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow\mathbb{Q}^{\mathfrak{A}}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{\mathbb{Q}^{\mathfrak{A}}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$. Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $\mathbb{Q}^{\mathfrak{A}}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $\mathbb{Q}^{\mathfrak{A}}$ is CSP-injective.
@@
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $\mathbb{Q}^{\mathfrak{A}}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $\mathbb{Q}^{\mathfrak{A}}$. By the CSP-injectivity of $\mathbb{Q}^{\mathfrak{A}}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $\mathbb{Q}^{\mathfrak{A}}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow\mathbb{Q}^{\mathfrak{A}}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{\mathbb{Q}^{\mathfrak{A}}}$ restricted to $i(C)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow\mathbb{Q}^{\mathfrak{A}}$.
@@
We now consider $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$. By Proposition 4.3, $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ is Pol-injective and $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}}_{I_{4}})$ and $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ are Datalog-interreducible. That $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}}_{I_{4}})$ and $\mathrm{CSP}(\mathfrak{A}$) are Datalog-interreducible then follows from the fact that Datalog-reductions can be composed.

### v2

@@ 3. Basic definitions
An embedding is an injective homomorphism $h\colon\mathfrak{A}\rightarrow\mathfrak{B}$ that additionally satisfies the following condition: for every $k$-ary relation symbol $R\in\tau$ and $\bar{t}\in A^{k}$ we have $h(\bar{t})\in R^{\mathfrak{B}}$ only if $\bar{t}\in R^{\mathfrak{A}}.$ An isomorphism is a surjective embedding. Two structures $\mathfrak{A}$ and $\mathfrak{B}$ are isomorphic if there exists an isomorphism from $\mathfrak{A}$ to $\mathfrak{B}$. An automorphism is an isomorphism from $\mathfrak{A}$ to $\mathfrak{A}$; we denote by $\mathrm{Aut}(\mathfrak{A})$ the set (group) of all automorphisms of $\mathfrak{A}$. The orbit of a tuple $\bar{t}\in A^{k}$ in $\mathfrak{A}$ is the set $\{g(\bar{t})\mid g\in\mathrm{Aut}(\mathfrak{A})\}$. Any tuples belonging to the same orbit satisfy precisely the same first-order formulas over $\mathfrak{A}$. The orbit equivalence relation provides a natural way to factorise relational structures; in the present paper, we will only need the following specific case. Given a relational structure $\mathfrak{A}$ and a subgroup $G\subseteq\mathrm{Aut}(\mathfrak{A})$, we denote by $\mathfrak{A}_{\!{}_{\vphantom{/}\smash{/G}}}$ the structure $\mathfrak{A}/E$ for
$E\coloneqq\{(x,y)\in A^{2}\mid\exists g\in G\colon g(x)=y\}.$
3.2. Universal algebra.
@@
An (equational) condition is a set of identities, i.e. formal expressions of the form $s\approx t$, where $s$ and $t$ are terms over a common set of function symbols. An equational condition is height-1 if it contains neither nested terms nor terms consisting of a single variable. We say that $\mathrm{Pol}(\mathfrak{A})$ (or some set of operations on $A$) satisfies an equational condition $\Sigma$ if the function symbols can be interpreted as elements of $\mathrm{Pol}(\mathfrak{A})$ so that, for each identity $s\approx t$ in $\Sigma$, the equality $s=t$ holds for any evaluation of variables in $A$. An example is the cyclic identity of arity $n\geq 2$ given by $f(x_{1},\dots,x_{n})\approx f(x_{2},\dots,x_{n},x_{1})$, for a symbol $f$ of arity $n\geq 2$. We say that an operation is cyclic if it satisfies the cyclic identity.
@@ 4.1. Injectivity and Datalog reductions
In the context of Conjecture 1.4, the concepts of CSP-injectivity and having no algebraicity are closely related. On the one hand, if $\mathfrak{B}$ is a homogeneous structure with no algebraicity and its relations only contain tuples with pairwise distinct entries, then it is CSP-injective [15, Lemma 4.3.6]. On the other hand, we can prove the following.
@@
We shall now see that CSP-injectivity and Pol-injectivity cannot hold simultaneously in $\omega$-categorical structures; yet, we shall observe immediately thereafter that CSP-injectivity can be traded against Pol-injectivity using the relation $I_{4}$ from Example 2.1.
@@
The expansion $\mathfrak{A}_{I_{4}}$ of any CSP-injective structure $\mathfrak{A}$ by the relation $I_{4}$ is Pol-injective, and $\mathrm{CSP}(\mathfrak{A})$ and $\mathrm{CSP}(\mathfrak{A}_{I_{4}})$ are Datalog-interreducible.
@@
Datalog reductions between CSPs are typically specified using factorisation-free Datalog-interpretations with parameters [1, 38], which are a particular case of logical interpretations with parameters. In the present paper, we only need a very basic understanding of Datalog-reductions, which can be easily explained on an informal level. Datalog is defined by adding formation rules to the existential positive fragment of first-order logic whose semantics is defined with inflationary fixed-points of definable operators. Every existential positive first-order formula is a Datalog formula and, if $\phi(\bar{x})$ is a Datalog formula over some relational signature $\tau\cup\{R\}$ with $R\notin\tau$, then $[\mathsf{ifp}_{R}\,\phi](\bar{x})$ is a Datalog formula over the signature $\tau$ whose semantics is given as follows. For a $\tau$-structure $\mathfrak{A}$ and a tuple $\bar{a}$ over $A$ matching the arity of $\bar{x}$, say $k$, we have $\mathfrak{A}\models[\mathsf{ifp}_{R}\,\phi](\bar{a})$ if and only if $\bar{a}$ is contained in the inflationary fixed-point of the operator $F^{\mathfrak{A}}_{\phi,R}(X)\coloneqq\{\bar{x}\in A^{k}\mid\mathfrak{A}_{X}\models\phi(\bar{x})\}$, i.e., the limit of the sequence $X_{0}\coloneqq\emptyset$ and $\smash{X_{i+1}\coloneqq X_{i}\cup F^{\mathfrak{A}}_{\phi,R}(X_{i})}$. For example, the Datalog formula
$\phi_{<}^{tc}(x,z)\coloneqq[\mathsf{ifp}_{T}\,(x<z)\vee(\exists y\ldotp T(x,y)\wedge T(y,z))](x,z)$
computes the transitive closure of $<$.
@@
For a relational $\tau$-structure $\mathfrak{A}$, we say that $\mathrm{CSP}(\mathfrak{A})$ is solvable in Datalog if there exists a Datalog sentence defining the class of all finite $\tau$-structures which do not homomorphically map to $\mathfrak{A}$. For example, $\mathrm{CSP}(\mathbb{Q};<)$ is solvable in Datalog using the sentence $\exists u\ldotp\phi_{<}^{tc}(u,u)$. It is not hard to see that every Datalog formula is specified by a finite set of rules of the form $R(\bar{x})\Leftarrow R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{m}(\bar{x}_{m})$ where $R$ is a fixed-point variable. In the case of the above Datalog formula computing the transitive closure of $<$, for example, the rules are $T(x,z)\Leftarrow(x<z)$ and $T(x,z)\Leftarrow T(x,y)\wedge T(y,z)$. When it comes to Datalog-reductions between CSPs, we only need the following intuitive understanding of this concept.
@@
A Datalog-interpretation is a mapping $\mathcal{I}$ from finite $\sigma$-structures to finite $\tau$-structures such that for every finite $\sigma$-structure $\mathfrak{J}$ the $\tau$-structure $\mathcal{I}(\mathfrak{J})$ can be defined from $\mathfrak{J}$ or a fixed finite power of it using Datalog formulas [1, Definition 1]. For a $\tau$-structure $\mathfrak{A}$ and $\sigma$-structure $\mathfrak{A^{\prime}}$, we say that $\mathrm{CSP}(\mathfrak{A^{\prime}})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$ if there exists a Datalog interpretation $\mathcal{I}$ so that $\mathfrak{J}\rightarrow\mathfrak{A^{\prime}}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow\mathfrak{A}$. Datalog reductions compose because Datalog interpretations do, i.e., if $\mathrm{CSP}(\mathfrak{A}^{\prime\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A}^{\prime})$ and $\mathrm{CSP}(\mathfrak{A}^{\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$, then $\mathrm{CSP}(\mathfrak{A}^{\prime\prime})$ Datalog-reduces to $\mathrm{CSP}(\mathfrak{A})$.
@@ 4.2. Wreath products
Let $\mathfrak{A}$ and $\mathfrak{B}$ be structures over disjoint relational signatures $\tau$ and $\sigma$. Moreover, let $E$ be a fresh binary symbol. The wreath product $\mathfrak{A}\wr\mathfrak{B}$ is the structure over the signature $\tau\cup\{E\}\cup\sigma$ with domain $A\times B$ whose relations are defined as follows. First,
$E^{\mathfrak{A}\wr\mathfrak{B}}\coloneqq\{((x_{1},y_{1}),(x_{2},y_{2}))\in(A\times B)^{2}\mid y_{1}=y_{2}\}.$
Then, for every $R\in\tau$ of arity $k$:
$R^{\mathfrak{A}\wr\mathfrak{B}}\coloneqq\{((x_{1},y_{1}),\dots,(x_{k},y_{k}))\in(A\times B)^{k}\mid(x_{1},\dots,x_{k})\in R^{\mathfrak{A}}\text{ and }y_{1}=\cdots=y_{n}\},$
and, for every $S\in\sigma$ of arity $m$:
$S^{\mathfrak{A}\wr\mathfrak{B}}\coloneqq\{((x_{1},y_{1}),\dots,(x_{m},y_{m}))\in(A\times B)^{m}\mid(y_{1},\dots,y_{m})\in S^{\mathfrak{B}}\}.$
@@
Given groups $G$ and $H$ acting on sets $A$ and $B$, respectively, their wreath product $G\wr H$ is given by $G^{B}\times H$ acting on $A\times B$ via $\left((g_{b})_{b\in B},h)\right)(a,b)\coloneqq(g_{h(b)}(a),h(b)).$ The wreath product $G\wr H$ again forms a group, with the neutral element $((e_{G})_{b\in B},e_{H})$ and the group operation
$((g_{b})_{b\in B},h))\cdot((g^{\prime}_{b})_{b\in B},h^{\prime}))\coloneqq((g_{b}\cdot g^{\prime}_{h^{-1}(b)})_{b\in B},h\cdot h^{\prime}))$.
Note that the permutation group on $A\times B$ induced by the component-wise action of the Cartesian product $G\times H$ is contained by that induced by the action of $G\wr H$ thereon: the permutations induced stemming from the former are represented in $G\wr H$ by all elements of the form $((g)_{b\in B},h)$.
@@
Item 6 in Proposition 4.5 follows from a famous theorem of Kechris, Pestov, and Todorčević [52] stating that a homogeneous structure $\mathfrak{B}$ with a countable relational signature is Ramsey if and only if its automorphism group $\mathrm{Aut}(\mathfrak{B})$ is extremely amenable, and hence the Ramsey property only depends on $\mathrm{Aut}(\mathfrak{B})$ viewed as an abstract topological group. Kechris, Pestov, and Todorčević observed that extreme amenability is preserved under taking wreath products; although they do not state this result explicitly, they do provide an example as a proof of concept: the automorphism group of the random convexly ordered equivalence relation (Example 4.4). Scow [64, Theorem 5.13] does state this result, and essentially provides a full proof, modulo Theorem 4.7 in [52].
4.3. Removing algebraicity
@@
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. As shown in Examples 1.3 and 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It then follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$ from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in ×$\mathbb{Q}\times\{a\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
@@
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
We now consider $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. By Proposition 4.3, $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is Pol-injective and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ are Datalog-interreducible. That $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}(\mathfrak{A}$) are Datalog-interreducible then follows from the fact that Datalog-reductions can be composed.